Michelle Walker is saving to buy a house in five years. She plans to put 20 percent down at that time, and she believes that she will need $28,000 for the down payment. If Michelle can invest in a fund that pays 6.80 percent annual interest, compounded quarterly, how much will she have to invest today to have enough money for the down payment?
Answer #1

We use the formula:
A=P(1+r/400)^4n
where
A=future value
P=present value
r=rate of interest
n=time period.

28000=P(1+0.068/4)^(4*5)

P=28000/(1+0.068/4)^(4*5)

=$28000*0.713807228

=$19986.60(Approx).